### v2

# Existence of solution to a system of PDEs modeling the crystal growth inside lithium batteries

Existence of solution to a system of PDEs modeling the crystal growth inside lithium batteries
Alexandros Skouras Institut Montpelliérain Alexander Grothendieck, Université de Montpellier, France alskouras@gmail.com Omar Lakkis Department of Mathematics, University of Sussex, Brighton UK lakkis.o.math@gmail.com  and  Vanessa Styles Department of Mathematics, University of Sussex, Brighton UK v.styles@sussex.ac.uk
(Date: March 18, 2025)
[…]
We study a model for lithium (Li) electrodeposition on Li-metal electrodes that leads to dendritic pattern formation. The model comprises of a system of three coupled PDEs, taking the form of an Allen–Cahn equation, a Nernst–Planck equation and a Poisson equation. We prove existence of a weak solution and stability results for this system and present numerical simulations resulting from a finite element approximation of the system, which illustrate the dendritic nature of solutions to the model.
1. Introduction
In this paper we analyse a model for lithium (Li) electrodeposition on Li-metal electrodes in high-capacity lithium-oxygen and lithium-sulphur batteries. The model we consider is a variant of the thermodynamically consistent model introduced in Chen et al. (2015) to predict dendritic patterns which form during an electrochemical process using Li-electrodeposition. For ease of presentation we consider a simplified version in which, where possible, physical parameters have been set to $1$:
$\partial_{t}u=\nabla\cdot[\boldsymbol{{A}}(\nabla u)\nabla u]-g^{\prime}(u)+m(c)h^{\prime}(u)\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{ in }\Omega\times[0,T],$ (1.1)
$\partial_{t}c=\nabla\cdot\big{[}D(u)\nabla c+D_{1}(u,c)\nabla\phi\big{]}-\partial_{t}u\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{ in }\Omega\times[0,T],$ (1.2)
$\nabla\cdot\big{[}\sigma(u)\nabla\phi\big{]}=\partial_{t}u-\sigma^{\prime}(u)\partial_{x}u\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{ in }\Omega\times[0,T],$ (1.3)
[…]
As humanity aims at reaching “net-zero” by 2050, replacing fossil fuels with renewable alternatives has spurred research in storage technologies, with lithium batteries being one of the favoured technologies. Electrodeposition in lithium batteries is an important phenomena that may lead to solid dendritic structures which, whilst remaining attached to the electrode, grow into the electrolyte solution. These structures can seriously compromise the performance and safety of such batteries, see for example Okajima et al. (2010); Akolkar (2013); Liang and Chen (2014); Chen et al. (2015); Yurkiv et al. (2018) and the references therein.
[…]
In this paper we adapt techniques from Burman and Rappaz (2003) in which the authors establish an existence result for weak solutions to a simplified version of the model (1.1) - (1.3). Specifically, the model studied in Burman and Rappaz (2003), which does not include an equation of type (1.3), is a system of two equations that take similar forms to (1.1) and (1.2). There are two key differences between the forms of (1.1) and (1.2) in this paper and their respective forms in Burman and Rappaz (2003), firstly, in this work the $\partial_{t}u$ term in (1.2) is not included in Burman and Rappaz (2003) and secondly, the $\nabla\phi$ advection term in (1.2) in this work is replaced by a $\nabla u$ advection term in Burman and Rappaz (2003).
The remainder of the paper is organised as follows. In Section 2 we present the full model together with a weak formulation, then we present the main result, which is an existence result, using Rothe’s method, for solutions to the weak formulation. We prove the existence result by considering, in Section 3, a time discretization of the weak formulation and, in Section 4, taking the limit as the time step $\tau$, of the dicretization, tends to zero. We conclude with Section 5 in which we present a finite element discretisation of the model together with some numerical examples that illustrate the dendritic nature of solutions to the model.
[…]
2. Weak formulation of the model
2.1. The model
The model, (1.1)–(1.3), takes the form of three interacting PDEs:
  1. (1.1)
    a nonlinear anisotropic Allen–Cahn type equation, for a phase-field variable $u$, with a forcing term, $m(c)h^{\prime}(u)$, that accounts for the Butler–Volmer electrochemical reaction kinetics, which model the concentration of lithium atoms;
  2. (1.2)
    a reaction–diffusion Nernst–Planck type equation, which describes the dynamics of the concentration, $c$, of lithium ions;
  3. (1.3)
    a Poisson type equation that describes the electro static potential, $\phi$, that drives the dynamics of the lithium ions.
For completeness we now rewrite the model together with the boundary and initial data:
$\partial_{t}u=\nabla\cdot[\boldsymbol{{A}}(\nabla u)\nabla u]-g^{\prime}(u)+m(c)h^{\prime}(u)\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{ in }\Omega\times[0,T],$ (2.1)
$\partial_{t}c=\nabla\cdot\big{[}D(u)\nabla c+D_{1}(u,c)\nabla\phi\big{]}-\partial_{t}u\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{ in }\Omega\times[0,T],$ (2.2)
$\nabla\cdot\big{[}\sigma(u)\nabla\phi\big{]}=\partial_{t}u-\sigma^{\prime}(u)\partial_{x}u\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{ in }\Omega\times[0,T],$ (2.3)
$\boldsymbol{{A}}(\nabla u)\nabla u\cdot\mathbf{n}=0,\leavevmode\nobreak\ \leavevmode\nobreak\ D(u)\nabla c\cdot\mathbf{n}=D_{1}(u,c)\nabla\phi\cdot\mathbf{n}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{ on }\partial\Omega\times[0,T],$ (2.4)
$\sigma(u)\nabla\phi\cdot\mathbf{n}=0\text{ on }\Gamma_{3}\times[0,T],\leavevmode\nobreak\ \phi=\phi_{-}\text{ on }\Gamma_{1}\times[0,T],\leavevmode\nobreak\ \phi=0\text{ on }\Gamma_{2}\times[0,T],$ (2.5)
$u(0)=u_{0},\leavevmode\nobreak\ \leavevmode\nobreak\ c(0)=c_{0}\text{ in }\Omega.$ (2.6)
Here $[0,T]$ is the time interval, the domain $\Omega=[0,L_{1}]\times[0,L_{2}]^{d-1}\subset\mathbb{R}^{d}$, $d=2,3$, is bounded and convex with Lipschitz boundary $\partial\Omega=\Gamma_{1}\cup\Gamma_{2}\cup\Gamma_{3}$, such that
$\Gamma_{1}:=\left\{\boldsymbol{x}\in\partial\Omega:\;x_{1}=0\right\},\leavevmode\nobreak\ \Gamma_{2}:=\left\{\boldsymbol{x}\in\partial\Omega:\;x_{1}=L_{1}\right\}\text{ and }\Gamma_{3}=\partial\Omega\smallsetminus(\Gamma_{1}\cup\Gamma_{2}),$ (2.7)
$g(u):=u^{2}(1-u)^{2}/4,\leavevmode\nobreak\ h(u):=u^{3}(6u^{2}-15u+10),$ (2.8)
$D(u):=D^{\text{e}}h(u)+D^{\text{s}}(1-h(u)),\leavevmode\nobreak\ \leavevmode\nobreak\ \sigma(u)=\sigma^{\text{e}}h(u)+\sigma^{\text{s}}(1-h(u)),\leavevmode\nobreak\ \text{ for }u\in\mathbb{R},$ (2.9)
$m(c):=\left\{\begin{array}[]{ll}C_{1}-C_{2},&c>1,\\ C_{1}-cC_{2},&0\leq c\leq 1,\\ C_{1},&c<0,\end{array}\right.\leavevmode\nobreak\ \leavevmode\nobreak\ D_{1}(u,c):=\left\{\begin{array}[]{ll}D(u),&c>1,\\ D(u)c,&0\leq c\leq 1,\\ 0,&c<0,\end{array}\right.$ (2.10)
with $\operatorname{D}\!^{e},D^{s},\sigma^{e},\sigma^{s}\in\mathbb{R}$ and $0<D^{\text{s}}<D^{\text{e}}$, $0<\sigma^{\text{e}}<\sigma^{\text{s}}$. The anisotropic diffusion tensor-valued field $\boldsymbol{{A}}(\boldsymbol{p})$ of $\boldsymbol{p}\in\mathbb{R}^{d}\smallsetminus{\left\{{\boldsymbol{0}}\right\}}$ is defined in Section 2.4.
2.2. Remark (cut-off of constitutive functions)
In Chen et al. (2015) the constitutive functions $D_{1}(u,c)$ and $m(c)$ are defined, for all $u,c\in\mathbb{R}$, to be $D_{1}(u,c)=D(u)c$ and $m(c)=C_{1}-cC_{2}$ respectively. However to establish the analytical results in this paper, we consider their cut off variants defined in 2.10.
2.3. The weak formulation
A weak formulation of (2.1)-(2.6) is given as follows:
Let $u(0)=u_{0}\in\operatorname{H}^{1}(\Omega)$ and $c(0)=c_{0}\in\operatorname{L}_{2}(\Omega)$. Then, find $(u,c,\phi)$
such that for all $(\chi,\eta)\in\operatorname{H}^{1}(\Omega)\times\operatorname{H}^{1}_{0|\Gamma_{1}\cup\Gamma_{2}}(\Omega)$,
$(\partial_{t}u,\chi)+(\boldsymbol{{A}}(\nabla u)\nabla u,\nabla\chi)+(g^{\prime}(u),\chi)=(m(c)h^{\prime}(u),\chi),$ (2.11a)
$(\partial_{t}c,\chi)+(D(u)\nabla c,\nabla\chi)+(D_{1}(u,c)\nabla\phi,\nabla\chi=(\partial_{t}u,\chi),$ (2.11b)
$(\sigma(u)\nabla\phi,\nabla\eta)=(-\partial_{t}u+\sigma^{\prime}(u)\partial_{x}u,\eta),$ (2.11c)
where
$g^{\prime}(s)=s(s-1)(s-\frac{1}{2}),\leavevmode\nobreak\ h^{\prime}(s)=30s^{2}(s-1)^{2}$ (2.12)
and
[…]
2.4. Anisotropic diffusion tensor
We define the tensor-valued field $\boldsymbol{{A}}(\boldsymbol{p})$, $\boldsymbol{p}\in\mathbb{R}^{d}\smallsetminus{\left\{{\boldsymbol{0}}\right\}}$ in (1.1), which models the anisotropy of the growing crystals due to the cubic crystalline structure of lithium, by
$\boldsymbol{{A}}(\boldsymbol{p})=a(\boldsymbol{p})\nabla a(\boldsymbol{p}){\boldsymbol{p}}^{\intercal}+a^{2}(\boldsymbol{p})\boldsymbol{\mathrm{I}}{}\\ =16\,\delta\,a(\boldsymbol{p}){\left|\boldsymbol{p}\right|}^{-6}\left[{p_{i}p_{j}({p_{i}}^{2}{\left|\boldsymbol{p}\right|}^{2}-\sum_{k=1}^{d}{p_{k}}^{4})}\right]_{i={1},\dotsc,{d}}^{j={1},\dotsc,{d}}+a^{2}(\boldsymbol{p})\boldsymbol{\mathrm{I}}{},$ (2.14)
where
$a(\boldsymbol{p})=a_{0}(1-3\delta)\left({1+\frac{4\delta}{1-3\delta}\frac{\sum_{j=1}^{d}{p_{j}}^{4}}{{\left|\boldsymbol{p}\right|}^{4}}}\right),$ (2.15)
for some material dependent parameters $a_{0},\delta>0$. Here $\boldsymbol{{A}}(\boldsymbol{p})$, $\boldsymbol{p}\in\mathbb{R}^{d}\smallsetminus{\left\{{\boldsymbol{0}}\right\}}$ arises from the (Fréchet) derivative, $J_{a}^{\prime}$, of the anisotropic Dirichlet energy,
$J_{a}(w):=\frac{1}{2}\int_{\Omega}a^{2}(\nabla w)|\nabla w|^{2}$ (2.16)
such that $J_{a}^{\prime}:\operatorname{H}^{1}(\Omega)\to(\operatorname{H}^{1}(\Omega))^{\prime},\leavevmode\nobreak\ \forall v\in\operatorname{H}^{1}(\Omega)$, satisfies
$J_{a}^{\prime}(w)[v]=\int_{\Omega}\boldsymbol{{A}}(\nabla w)\nabla w\cdot\nabla v,\leavevmode\nobreak\ \leavevmode\nobreak\ \forall v\in\operatorname{H}^{1}(\Omega).$ (2.17)
2.5. Lemma (convexity of the anisotropic Dirichlet energy)
Recalling (2.15), for $\delta<1/15$ , the functional
$j_{a}(\boldsymbol{p}):=\frac{1}{2}\int_{\Omega}a^{2}(\boldsymbol{p})|\boldsymbol{p}|^{2}$ (2.18)
is strictly convex in $\boldsymbol{p}$ for all $\boldsymbol{p}\in(\operatorname{L}_{2}(\Omega))^{d}$. It follows that the anisotropic Dirichlet energy $J_{a}(w)$ defined in (2.16) is strictly convex in $w\in\operatorname{H}^{1}(\Omega)$.
Chứng minh.
The result is proved in Burman and Rappaz (2003) for $d=2$. Here we extend it to the case $d=3$. A sufficient condition for a functional to be strictly convex is the positivity of the Hessian matrix, Dacorogna (2008), i.e. we require
$\mathcal{H}(\boldsymbol{p})=\nabla^{2}\bar{a}^{2}(\boldsymbol{p})>0,\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\boldsymbol{p}\in\operatorname{Dom}\bar{a}^{2},$
where $\bar{a}:\mathbb{R}^{d}\to\mathbb{R}$ is given by
$\bar{a}(\boldsymbol{p}):=\frac{\sqrt{2}a(\boldsymbol{p})}{2}|\boldsymbol{p}|$ (2.19)
such that
$j_{a}(\boldsymbol{p})=\frac{1}{2}\int_{\Omega}a^{2}(\boldsymbol{p})|\boldsymbol{p}|^{2}=\int_{\Omega}\bar{a}^{2}(\boldsymbol{p}).$ (2.20)
[…]
With the use of software Wolfram Mathematica Wolfram Research, Inc. (2024), our calculations show that $\mathcal{H}_{11}$ is positive if and only if
[…]
The above inequalities hold for $\delta<1/15$, while the determinant of the upper $2\times 2$ matrix is positive if and only if
[…]
which are statisfied if and only if $\delta<1/15$. Lastly, for the determinant of the Hessian matrix we have
$\det(\mathcal{H}(\boldsymbol{p}))>0\Leftrightarrow 1+6\delta-315\delta^{2}>0\text{ and }1+18\delta-15\delta^{2}>0,$ (2.23)
which also holds for $\delta<1/15$.
The convexity of $J_{a}$ follows from restricting $j_{a}$ to the linear subspace of $\operatorname{L}_{2}(\varOmega)$ comprising all gradients of $\operatorname{H}^{1}(\Omega)$ functions. ∎
2.6. Remark (monotonicity)
Since the anisotropic Dirichlet energy (2.16) is Fréchet differentiable and strictly convex it follows (from Pro 7.4 Showalter, 2013, e.g.) that the mapping $w\in\operatorname{H}^{1}(\Omega)\mapsto J_{a}^{\prime}(w)\in(\operatorname{H}^{1}(\Omega))^{\prime}$ is monotone and hemicontinuous. Thus we have
$J_{a}(w)-J_{a}(v)\leq J_{a}^{\prime}(w)[w-v]=\int_{\Omega}\boldsymbol{{A}}(\nabla w)\nabla w\nabla(w-v).$ (2.24)
2.7. Remark (ellipticity)
Noting Gräser (2011, Lem 2.2) we have that
$(\boldsymbol{{A}}(\boldsymbol{p})\boldsymbol{p},\boldsymbol{p})=(\nabla(\bar{a}^{2}(\boldsymbol{p})),\boldsymbol{p})=2\int_{\Omega}\bar{a}^{2}(\boldsymbol{p}),\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\boldsymbol{p}\in\mathbb{R}^{d}$ (2.25)
thus from (2.15) and (2.19) we establish the bounds
$c_{A}|w|^{2}_{\operatorname{H}^{1}(\Omega)}\leq\int_{\Omega}\boldsymbol{{A}}(\nabla w)|\nabla w|^{2}\leq C_{A}|w|^{2}_{\operatorname{H}^{1}(\Omega)}$ (2.26)
with
$c_{A}:=\min_{\boldsymbol{p}\in\mathbb{R}^{d}}a^{2}(\boldsymbol{p})\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ C_{A}:=\max_{\boldsymbol{p}\in\mathbb{R}^{d}}a^{2}(\boldsymbol{p}).$ (2.27)
2.8. Theorem (existence of solutions)
For $\delta<1/15$, recall (2.15), initial data $(u_{0},c_{0})\in\operatorname{H}^{1}(\Omega)\times\operatorname{L}_{2}(\Omega)$ and $T>0$, there exist
$u\in\operatorname{L}_{2}([0,T];\operatorname{H}^{1}(\Omega))\cap\operatorname{H}^{1}([0,T];\operatorname{L}_{2}(\Omega))\cap\operatorname{L}_{\infty}([0,T];\operatorname{L}_{\infty}(\Omega)),$
$c\in\operatorname{L}_{2}([0,T];\operatorname{H}^{1}(\Omega))\cap\operatorname{H}^{1}([0,T];(\operatorname{H}^{1}(\Omega))^{\prime})\cap\operatorname{L}_{\infty}([0,T];\operatorname{L}_{2}(\Omega))$
and
$\phi\in\operatorname{L}_{2}([0,T];\operatorname{H}^{1}_{0|\Gamma_{1}\cup\Gamma_{2}}(\Omega))$
such that $(u,c,\phi)$ satisfies $u(0)=u_{0}$, $c(0)=c_{0}$ and (2.11a)–(2.11c) for almost every $t\in[0,T]$ and every $(\chi,\eta)\in\operatorname{H}^{1}(\Omega)\times\operatorname{H}^{1}_{0|\Gamma_{1}\cup\Gamma_{2}}(\Omega)$.
We split the proof of Theorem 2.8 in two parts. The first part, which is presented in Section 3, consists of proving existence of a solution to a semi-implicit Euler time discretization of (2.11a)–(2.11c) and deriving stability bounds on this solution. In the second part, presented in Section 4, we show that in the limit as the time step $\tau$ of the semi-discrete approximation tends to zero, there exists a solution that solves the weak formulation (2.11a)–(2.11c).
3. Time discretization
In this section we consider a time discretization of (2.11a)–(2.11c). To this end, we partition the time interval $[0,T]$ into $K$ equidistant steps, $0=t_{0}<t_{1}<\ldots<t_{K-1}<t_{K}=T$, so that for $k=1,\ldots,K$, $t_{k}=k\tau$.
We use the backward difference quotient, $d_{\tau}w^{k}=(w^{k}-w^{k-1})/\tau$ to obtain the following discretization of (2.11a)–(2.11c) in which we have linearized the convection term in (2.11b) and derived a specific form of the right hand side of (2.11a) that lends itself to proving the boundedness of $u^{k}$ (see Lemma 3.2).
For all $(\chi,\eta)\in\operatorname{H}^{1}(\Omega)\times\operatorname{H}^{1}_{0|\Gamma_{1}\cup\Gamma_{2}}(\Omega)$ and for all $k\geq 1,k\in\mathbb{N}$
$(d_{\tau}u^{k},\chi)+(\boldsymbol{{A}}(\nabla u^{k})\nabla u^{k},\nabla\chi)+(g^{\prime}(u^{k}),\eta)=((m(c^{k-1}))^{-}h_{1}^{\prime}(u^{k}),\chi)$
$+((m(c^{k-1}))^{+}h_{2}^{\prime}(u^{k}),\chi),$ (3.1a)
$(\sigma(u^{k})\nabla\phi^{k},\nabla\eta)=(-d_{\tau}u^{k}+\sigma^{\prime}(u^{k})\partial_{x}u^{k},\eta),$ (3.1b)
$(d_{\tau}c^{k},\chi)+(D(u^{k})\nabla c^{k},\nabla\chi)=-(d_{\tau}u^{k},\chi)-(D_{1}(u^{k},c^{k-1})\nabla\phi^{k},\nabla\chi),$ (3.1c)
with $u^{0}=u(0)=u_{0}$, $c^{0}=c(0)=c_{0}$,
$h_{1}^{\prime}(u^{k}):=30u^{k-1}u^{k}(u^{k}-1)^{2}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ h_{2}^{\prime}(u^{k}):=30(u^{k})^{2}(u^{k}-1)(u^{k-1}-1)$ (3.2)
and $(\xi)^{-}:=\min(0,\xi)$ and $(\xi)^{+}:=\max(0,\xi)$.
Note that the equations in the system (3.1)–(3.1c) are not simultaneous, but will be solved in sequential order, first the phase-field equation (3.1) is solved with given $c^{k-1}$ to obtain $u^{k}$, then (3.1b) is solved with given $u^{k}$ to obtain $\phi^{k}$ and finally the convection-diffusion equation (3.1c) is solved with given $u^{k}$ and $\phi^{k}$ to obtain $c^{k}$.
We make the following assumptions on the time step $\tau$, the anisotropy parameter $\delta$, defined in (2.15), and the initial data $u_{0},c_{0}$,
$\begin{gathered}\displaystyle{0<\tau<\min\left(\nicefrac{{1}}{{2}},\nicefrac{{1}}{{C^{\star}}},\nicefrac{{1}}{{4C^{\ast}}}\right)},\\ \delta<\nicefrac{{1}}{{15}},\\ 0\leq u_{0}\leq 1,u_{0}\in\operatorname{H}^{1}(\Omega),c_{0}\in\operatorname{L}_{2}(\Omega).\end{gathered}$ (3.3)
Here $C^{\star}$ and $C^{\ast}$ are defined by the bounds in (3.23) and (3.27) respectively.
3.1. Theorem (existence and uniqueness of a time-discrete solution)
Let assumptions (3.3) hold, then there exists a unique triple
$(u^{k},c^{k},\phi^{k})\in(\operatorname{H}^{1}(\Omega))^{2}\times\operatorname{H}^{1}_{0|\Gamma_{1}\cup\Gamma_{2}}(\Omega),$ (3.4)
with $k=1,...,K=\lceil T/\tau\rceil$, satisfying the $k$-th step of (3.1)–(3.1c).
We split the proof of Theorem 3.1 into several parts consisting of lemmas 3.43.7, in which we prove, in turn, the existence of unique solutions, $u^{k}$, $\phi^{k}$ and $c^{k}$ to (3.1), (3.1b) and (3.1c) respectively. In order to show the existence of a unique solution, $u^{k}$ to (3.1), we first show that $u^{k}$ remains valued in $\left[0,1\right]$ for all $k$.
3.2. Lemma (stability of the time-discrete order parameter)
Under assumptions (3.3) and $0\leq u^{k-1}\leq 1$, we have
$0\leq u^{k}\leq 1$ (3.5)
Chứng minh.
Setting $\chi=(u^{k})^{-}$, in (3.1) gives
$(u^{k},(u^{k})^{-})+\tau(\boldsymbol{{A}}(\nabla u^{k})\nabla u^{k},\nabla(u^{k})^{-})-\tau((m(c^{k-1}))^{-}h_{1}^{\prime}(u^{k}),(u^{k})^{-})\\ -\tau((m(c^{k-1}))^{+}h_{2}^{\prime}(u^{k}),(u^{k})^{-})+\tau g^{\prime}(u^{k}),(u^{k})^{-})=(u^{k-1},(u^{k})^{-})$ (3.6)
then noting (2.26) we have
$\|(u^{k})^{-}\|_{\operatorname{L}_{2}(\Omega)}^{2}+\tau c_{A}|(u^{k})^{-}|^{2}_{\operatorname{H}^{1}(\Omega)}\leq(u^{k-1},(u^{k})^{-})+\tau((m(c^{k-1}))^{-}h_{1}^{\prime}(u^{k}),(u^{k})^{-})\\ +\tau((m(c^{k-1}))^{+}h_{2}^{\prime}(u^{k}),(u^{k})^{-})-\tau(g^{\prime}(u^{k}),(u^{k})^{-}).$ (3.7)
By assumption $(u^{k-1})^{-}=0$ and noting from (2.12) and (3.2) that $(g^{\prime}(u^{k}),(u^{k})^{-})\geq 0$, $((m(c^{k-1}))^{-}h_{1}^{\prime}(u^{k}),(u^{k})^{-})\leq 0$ and $((m(c^{k-1}))^{+}h_{2}^{\prime}(u^{k}),(u^{k})^{-})\leq 0$ it follows that
$\|(u^{k})^{-}\|^{2}_{\operatorname{L}_{2}(\Omega)}\leq 0\leavevmode\nobreak\ \leavevmode\nobreak\ \Rightarrow\leavevmode\nobreak\ \leavevmode\nobreak\ u^{k}\geq 0.$ (3.8)
Working similarly with $\eta=(u^{k}-1)^{+}$ we obtain $u^{k}\leq 1$, which concludes the proof. ∎
3.3. Remark (useful inequalities)
Combining the result in Lemma 3.2 with (2.9), (2.10) and (2.12), we have
$0<\min(D^{e},D^{s})=:D_{\min}\leq D(u^{k})\leq D_{\max}:=\max(D^{e},D^{s}),$ (3.9)
$0<\min(\sigma^{e},\sigma^{s})=:\sigma_{\min}\leq\sigma(u^{k})\leq\sigma_{\max}:=\max(\sigma^{e},\sigma^{s}),$ (3.10)
${\left|g^{\prime}(u)\right|}+{\left|h^{\prime}(u)\right|}+{\left|m(c)\right|}+{\left|D_{1}(u,c)\right|}\leq C,$ (3.11)
${\left|\left(g^{\prime}(u^{k}),\chi\right)\right|}\leq C\|u^{k}\|_{\operatorname{L}_{2}(\Omega)}\|\chi\|_{\operatorname{L}_{2}(\Omega)},$ (3.12)
${\left|\left((m(c^{k-1}))^{-}h_{1}^{\prime}(u^{k})+(m(c^{k-1}))^{+}h_{2}^{\prime}(u^{k}),\chi\right)\right|}\leq C\|u^{k}\|_{\operatorname{L}_{2}(\Omega)}\|\chi\|_{\operatorname{L}_{2}(\Omega)}.$ (3.13)
3.4. Lemma (existence and uniqueness of solutions for the time-discrete order parameter)
Suppose assumptions (3.3) hold and $u^{k-1}\in\operatorname{L}_{2}(\Omega)$. Then there exists a unique solution $u^{k}\in\operatorname{H}^{1}(\Omega)$ that satisfies the $k$-th step of (3.1).
Chứng minh.
We rewrite (3.1) as follows
$(d_{\tau}u^{k},\chi)+I^{\prime}(u^{k})\chi=0,\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\chi\in\operatorname{H}^{1}(\Omega),$ (3.14)
where $I^{\prime}:\operatorname{H}^{1}(\Omega)\to(\operatorname{H}^{1}(\Omega))^{\prime}$ is given by
$I^{\prime}(w)\chi=\left(\boldsymbol{{A}}(\nabla w)\nabla w,\nabla\chi\right)+\left(g^{\prime}(w)-(m(c^{k-1}))^{-}h_{1}^{\prime}(w)-(m(c^{k-1}))^{+}h_{2}^{\prime}(w),\chi\right)$ (3.15)
is the (Fréchet) derivative of the energy functional $I:\operatorname{H}^{1}(\Omega)\to\mathbb{R}$ with
$I(w)=\int_{\Omega}\Big{(}\frac{a(\nabla w)^{2}}{2}|\nabla w|^{2}+g^{\prime}(w)-(m(c^{k-1}))^{-}h_{1}^{\prime}(w)-(m(c^{k-1}))^{+}h_{2}^{\prime}(w)\Big{)}.$ (3.16)
We now seek a minimiser to the functional $I_{k}:\operatorname{H}^{1}(\Omega)\to\mathbb{R}$
$I_{k}(w)=\frac{1}{2\tau}\|w-u^{k-1}\|^{2}_{\operatorname{L}_{2}(\Omega)}+I(w),$ (3.17)
which is the functional from which we derive (3.14). From Theorem 3.30 in Dacorogna (2008), to show the existence of a minimiser of (3.17) it is enough to prove that the function $F:\Omega\times\mathbb{R}\times\mathbb{R}^{d}\to\mathbb{R}$, defined by
$F(x,\xi,\boldsymbol{p}):=\frac{a(\boldsymbol{p})^{2}}{2}|\boldsymbol{p}|^{2}+\frac{1}{2\tau}(\xi-u^{k-1})^{2}\\ +g^{\prime}(\xi)-(m(c^{k-1}))^{-}h_{1}^{\prime}(\xi)-(m(c^{k-1}))^{+}h_{2}^{\prime}(\xi),$ (3.18)
is a Carathéodory function, coercive and satisfies a sufficient growth condition. From (Bartels, 2015, Remark 2.4(i)) it follows that $F$ is a Carathédory function. From (3.11) we deduce that there are constants $\gamma_{1},\gamma_{2}>0$ such that
$-\gamma_{1}\leq g^{\prime}(\xi)-(m(c^{k-1}))^{-}h_{1}^{\prime}(\xi)-(m(c^{k-1}))^{+}h_{2}^{\prime}(\xi)\leq\gamma_{2}.$ (3.19)
Using (2.27) and (3.19) we establish the following bounds on $F$:
$F(x,\xi,\boldsymbol{p})$ $=\frac{a(\boldsymbol{p})^{2}}{2}|\boldsymbol{p}|^{2}+\frac{1}{2\tau}(\xi-u^{k-1})^{2}+g^{\prime}(\xi)-(m(c^{k-1}))^{-}h_{1}^{\prime}(\xi)$
$\qquad-(m(c^{k-1}))^{+}h_{2}^{\prime}(\xi)$
$\geq\frac{c_{A}}{2}{\left|\boldsymbol{p}\right|}^{2}+\frac{1}{4\tau}\xi^{2}-\frac{1}{2\tau}(u^{k-1})^{2}-\gamma_{1}$ (3.20)
and
$F(x,\xi,\boldsymbol{p})$ $=\frac{a(\boldsymbol{p})^{2}}{2}|\boldsymbol{p}|^{2}+\frac{1}{2\tau}(\xi-u^{k-1})^{2}+g^{\prime}(\xi)-(m(c^{k-1}))^{-}h_{1}^{\prime}(\xi)$
$\qquad-(m(c^{k-1}))^{+}h_{2}^{\prime}(\xi)$
$\leq\frac{C_{A}}{2}{\left|\boldsymbol{p}\right|}^{2}+\frac{1}{\tau}\xi^{2}+\frac{1}{\tau}(u^{k-1})^{2}+\gamma_{2}.$ (3.21)
It only remains to show the uniqueness of the minimizer. Assume $w_{1}$ and $w_{2}$ are minimizers. From (3.14) it follows that
$\int_{\Omega}\frac{1}{\tau}(w_{1}-w_{2})v+\int_{\Omega}(\boldsymbol{{A}}(\nabla w_{1})\nabla w_{1}-\boldsymbol{{A}}(\nabla w_{2})\nabla w_{2})\cdot\nabla v\\ +\,\int_{\Omega}\Big{(}g^{\prime}(w_{1})-g^{\prime}(w_{2})\Big{)}v-\int_{\Omega}(m(c^{k-1}))^{+}\Big{(}h_{1}^{\prime}(w_{1})-h_{1}^{\prime}(w_{2})\Big{)}v\\ -\int_{\Omega}(m(c^{k-1}))^{-}\Big{(}h_{2}^{\prime}(w_{1})-h_{2}^{\prime}(w_{2})\Big{)}v=0.$ (3.22)
Choosing $v=w_{1}-w_{2}$ in (3.22), using the Lipschitz continuity of $g^{\prime}$, $h_{1}^{\prime}(u^{k})$ and $h_{2}^{\prime}(u^{k})$, (3.11), and noting from Remark 2.6 the monotonicity of the mapping $w\mapsto J_{a}^{\prime}(w)$, we obtain
$\frac{1}{\tau}\|w_{1}-w_{2}$ $\|^{2}_{\operatorname{L}_{2}(\Omega)}\leq\int_{\Omega}|(g^{\prime}(w_{1})-g^{\prime}(w_{2}))(w_{1}-w_{2})|$ (3.23)
$+C\int_{\Omega}|h_{1}^{\prime}(w_{1})-h_{1}^{\prime}(w_{2})||w_{1}-w_{2}|\leq C^{\star}\|w_{1}-w_{2}\|^{2}_{\operatorname{L}_{2}(\Omega)}$
$+C\int_{\Omega}|h_{2}^{\prime}(w_{1})-h_{2}^{\prime}(w_{2})||w_{1}-w_{2}|\leq C^{\star}\|w_{1}-w_{2}\|^{2}_{\operatorname{L}_{2}(\Omega)}$
and hence
$(\frac{1}{\tau}-C^{\star})\|w_{1}-w_{2}\|^{2}_{\operatorname{L}_{2}(\Omega)}\leq 0,$ (3.24)
where $C^{\star}$ is independent of $\tau$. Thanks to assumption (3.3) we have $\tau<1/C^{\ast}$ and hence the uniqueness of the minimizer. ∎
3.5. Lemma (energy estimates for the order parameter)
Under assumptions (3.3) we have
$\max_{l=0,\dotsc,K}\|u^{l}\|_{\operatorname{L}_{2}(\Omega)}^{2}+\tau\sum_{k=1}^{K}\|u^{k}\|^{2}_{\operatorname{H}^{1}(\Omega)}+\tau\sum_{k=1}^{K}\|d_{\tau}u^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}\leq C.$ (3.25)
Chứng minh.
Setting $\chi=u^{k}$ in (3.14) and noting (2.26), (3.15), (3.13), (3.12) and
$(d_{\tau}u^{k},u^{k})=\frac{\tau}{2}\|d_{\tau}u^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}+\frac{1}{2}d_{\tau}\|u^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}$ (3.26)
we have
$\frac{\tau}{2}\|d_{\tau}u^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}\!\!\!$ $+$ $\!\!\!\frac{1}{2}d_{\tau}\|u^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}+c_{A}|u^{k}|^{2}_{\operatorname{H}^{1}(\Omega)}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ {\left|\left(g^{\prime}(u^{k}),u^{k}\right)\right|}$ (3.27)
$+{\left|\left((m(c^{k-1}))^{-}h_{1}^{\prime}(u^{k}),u^{k}\right)\right|}+{\left|\left((m(c^{k-1}))^{+}h_{2}^{\prime}(u^{k}),u^{k}\right)\right|}$
$\qquad\qquad\qquad\qquad\qquad\qquad\leq C^{\ast}\|u^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)},$
where $C^{\ast}$ is independent of $\tau$. Multiplying by $\tau$ and summing from $k=1,2,...,l$, yields for any $l\leq K$ that
$\left(\frac{1}{2}-\tau C^{\ast}\right)\|u^{l}\|^{2}_{\operatorname{L}_{2}(\Omega)}+\tau c_{A}\sum_{k=1}^{l}|u^{k}|^{2}_{\operatorname{H}^{1}(\Omega)}\leq\frac{1}{2}\|u^{0}\|^{2}_{\operatorname{L}_{2}(\Omega)}+\tau C\sum_{k=1}^{l-1}\|u^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}.$ (3.28)
Owing to (3.3) we have $\tau<1/(4C^{\star})$ and hence using a generalized discretized Gronwall lemma, see (Bartels, 2015, Lem 2.2), we have that
$\frac{1}{4}\max_{l={1},\dotsc,{K}}\|u^{l}\|^{2}_{\operatorname{L}_{2}(\Omega)}+\tau c_{A}\sum_{k=1}^{K}|u^{k}|_{\operatorname{H}^{1}(\Omega)}^{2}\leq\frac{1}{2}\|u^{0}\|^{2}_{\operatorname{L}_{2}(\Omega)}\operatorname{e}\!^{CT}\leq C,$ (3.29)
which yields the first two bounds in (3.25). To prove the third bound we set $\chi=d_{\tau}u^{k}$ in (3.14), multiply the resulting equation by $\tau$ and sum over $k=1,2,...,l$, to obtain for any $l\leq K$ that
$\tau\sum_{k=1}^{l}\|d_{\tau}u^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}+\sum_{k=1}^{l}(\boldsymbol{{A}}(\nabla u^{k})\nabla u^{k},\nabla[u^{k}-u^{k-1}])+\tau\sum_{k=1}^{l}(g^{\prime}(u^{k}),d_{\tau}u^{k})\\ \qquad=\tau\sum_{k=1}^{l}((m(c^{k-1}))^{-}h_{1}^{\prime}(u^{k}),d_{\tau}u^{k})+\tau\sum_{k=1}^{l}((m(c^{k-1}))^{+}h_{2}^{\prime}(u^{k}),d_{\tau}u^{k}).$ (3.30)
Using (2.24), (3.13) and (3.12) we have
$\tau\sum_{k=1}^{l}\|d_{\tau}u^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}+\sum_{k=1}^{l}(J_{a}(u^{k})-J_{a}(u^{k-1}))\leq C\tau\sum_{k=1}^{l}\|u^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}+\frac{\tau}{2}\sum_{k=1}^{l}\|d_{\tau}u^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}.$ (3.31)
Hence, noting (3.29), (3.3) and $u_{0}\in\operatorname{H}^{1}(\Omega)$, we have
$\frac{\tau}{2}\sum_{k=1}^{l}\|d_{\tau}u^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}+J_{a}(u^{l})\leq J_{a}(u^{0})+\frac{\tau C}{4}\sum_{k=1}^{l}\|u^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}\leq C,$ (3.32)
which completes the proof. ∎
3.6. Lemma (existence and uniqueness of the time-discrete potential)
Let assumptions (3.3) hold. Then there exists a unique solution $\phi^{k}\in\operatorname{H}^{1}_{0|\Gamma_{1}\cup\Gamma_{2}}(\Omega)$, that satisfies the $k$-th step of (3.1b).
Chứng minh.
We introduce the form $b_{1}:(\operatorname{H}^{1}(\Omega))^{2}\to\mathbb{R}$
$b_{1}(u^{k};v,w)=\int_{\Omega}\sigma(u^{k})\nabla v\cdot\nabla w,$ (3.33)
which is bilinear in respect to $v$ and $w$. Using (3.33) we can rewrite (3.1b) as
$b_{1}(u^{k};\phi^{k},\eta)=-(d_{\tau}u^{k}+\sigma^{\prime}(u^{k})\partial_{x}u^{k},\eta).$ (3.34)
We now prove existence of a weak solution using the Lax–Milgram Theorem. The coercivity and boundedness of $b_{1}$ follow from (3.10) and the equivalence of the $\operatorname{H}^{1}$ semi-norm and the $\operatorname{H}^{1}$ norm under a homogeneous Dirichlet boundary condition:
$b_{1}(u^{k};v,v)=\int_{\Omega}\sigma(u^{k}){\left|\nabla v\right|}^{2}\geq\sigma_{\min}{\left|v\right|}^{2}_{\operatorname{H}^{1}(\Omega)},$ (3.35)
${\left|b_{1}(u^{k};v,w)\right|}\leq\sigma_{\max}{\left|v\right|}_{\operatorname{H}^{1}(\Omega)}{\left|w\right|}_{\operatorname{H}^{1}(\Omega)}.$ (3.36)
From Lemma 3.5 we have that
$\|u^{k}\|^{2}_{\operatorname{H}^{1}(\Omega)}+\|d_{\tau}u^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}\leq\frac{C}{\tau}$ (3.37)
which together with (2.13), imply
$\|d_{\tau}u^{k}+\sigma^{\prime}(u^{k})\partial_{x}u^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}\leq\frac{C}{\tau}.$ (3.38)
Thus, by the Lax–Milgram Theorem, (3.1b) has a unique solution at the $k$-th step $\phi^{k}\in\operatorname{H}^{1}_{0|\Gamma_{1}\cup\Gamma_{2}}(\Omega)$, with
$\|\phi^{k}\|_{\operatorname{H}^{1}(\Omega)}\leq\frac{C}{\tau}.$ (3.39)
3.7. Lemma (existence and uniqueness of the time-discrete concentration)
Under assumptions (3.3), if $\|c^{k-1}\|_{\operatorname{L}_{2}(\Omega)}\leq C$, then there exists a unique solution $c^{k}\in\operatorname{H}^{1}(\Omega)$ that satisfies the $k$-th step of (3.1c).
Chứng minh.
We introduce the form $b_{2}:(\operatorname{H}^{1}(\Omega))^{2}\to\mathbb{R}$
$b_{2}(u^{k};w,v):=\int_{\Omega}D(u^{k})\nabla w\cdot\nabla v+\frac{1}{\tau}wv,$ (3.40)
which is bilinear for $w$ and $v$. Using (3.40) we can rewrite (3.1c) as
$b_{2}(u^{k};c^{k},\chi)=-(d_{\tau}u^{k},\chi)-(D_{1}(u^{k},c^{k-1})\nabla\phi^{k},\nabla\chi)+\frac{1}{\tau}(c^{k-1},\chi).$ (3.41)
Again we prove existence of a weak solution using the Lax–Milgram Theorem. Coercivity and boundedness of $b_{2}$ follow from (3.9):
$b_{2}(u^{k};w,w)=\int_{\Omega}D(u^{k}){\left|\nabla w\right|}^{2}+\frac{1}{\tau}{\left|w\right|}^{2}\geq D_{\min}|w|^{2}_{\operatorname{H}^{1}(\Omega)}+\frac{1}{\tau}\|w\|^{2}_{\operatorname{L}_{2}(\Omega)}\\ \geq D_{\min}\|w\|^{2}_{\operatorname{H}^{1}(\Omega)},$ (3.42)
${\left|b_{2}(u^{k};w,v)\right|}={\left|\int_{\Omega}D(u^{k})\nabla w\cdot\nabla v+\frac{1}{\tau}wv\right|}\leq\int_{\Omega}|D(u^{k})||\nabla w||\nabla v|+\frac{1}{\tau}|w||v|\\ \leq D_{\max}|w|_{\operatorname{H}^{1}(\Omega)}|v|_{\operatorname{H}^{1}(\Omega)}+\frac{1}{\tau}\|w\|_{\operatorname{L}_{2}(\Omega)}\|v\|_{\operatorname{L}_{2}(\Omega)}\leq\frac{1}{\tau}\|w\|_{\operatorname{H}^{1}(\Omega)}\|v\|_{\operatorname{H}^{1}(\Omega)}.$ (3.43)
From Lemma 3.5, the assumption that $\|c^{k-1}\|_{\operatorname{L}_{2}(\Omega)}\leq C$, (3.11) and (3.39) we have
$\|d_{\tau}u^{k}\|_{\operatorname{L}_{2}(\Omega)}+\frac{1}{\tau}\|c^{k-1}\|_{\operatorname{L}_{2}(\Omega)}+\|D_{1}(u^{k},c^{k-1})\nabla\phi^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}\leq\frac{C}{\tau}.$ (3.44)
Thus the right hand side of (3.41) is bounded, dependent on $\tau$, in $(\operatorname{H}^{1}(\Omega))^{\prime}$ and by the Lax–Milgram Theorem, there exists a unique solution $c^{k}\in\operatorname{H}^{1}(\Omega)$ that satisfies the $k$-th step of (3.1c). ∎
3.8. Lemma (energy stability of the potential)
Let assumptions (3.3) hold and let $\phi^{k}$, be the solution to the $k$-th step of (3.1b). Then we have
$\tau\sum_{k=1}^{K}\|\phi^{k}\|^{2}_{\operatorname{H}^{1}(\Omega)}\leq C.$ (3.45)
Chứng minh.
We set $\eta=\phi^{k}$ in (3.1b) to give
$\int_{\Omega}\sigma(u^{k})|\nabla\phi^{k}|^{2}=-\int_{\Omega}d_{\tau}u^{k}\phi^{k}+\int_{\Omega}\sigma^{\prime}(u^{k})\partial_{x}u^{k}\phi^{k}.$
Noting (3.10) we have
$\leavevmode\nobreak\ \sigma_{\min}|\phi^{k}|^{2}_{\operatorname{H}^{1}(\Omega)}\leq\|d_{\tau}u^{k}\|_{\operatorname{L}_{2}(\Omega)}\|\phi^{k}\|_{\operatorname{L}_{2}(\Omega)}+\sigma^{\prime}_{\max}\|\partial_{x}u^{k}\|_{\operatorname{L}_{2}(\Omega)}\|\phi^{k}\|_{\operatorname{L}_{2}(\Omega)},$ (3.46)
where $\sigma_{\max}^{\prime}:=\max_{s\in\mathbb{R}}\sigma^{\prime}(s)$. Hence from the Friedrichs inequality we have
$\begin{split}\frac{\sigma_{\min}}{2C_{f}}\|\phi^{k}\|_{\operatorname{L}_{2}(\Omega)}^{2}+\frac{\sigma_{\min}}{2}|\phi^{k}|^{2}_{\operatorname{H}^{1}(\Omega)}\leq C\|d_{\tau}u^{k}\|_{\operatorname{L}_{2}(\Omega)}^{2}+\frac{\sigma_{\min}}{8C_{f}}\|\phi^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}\\ +C\|\partial_{x}u^{k}\|_{\operatorname{L}_{2}(\Omega)}^{2}+\frac{\sigma_{\min}}{8C_{f}}\|\phi^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}\end{split}$ (3.47)
where $C_{f}$ is the constant in the Friedrichs inequality, which is associated with the diameter of $\Omega$. After rearranging (3.47), multiplying the resulting inequality by $\tau$ and summing over $k=1,2,...,l$, for any $l\leq K$ we have
$\tau\sum_{k=1}^{l}\|\phi^{k}\|^{2}_{\operatorname{H}^{1}(\Omega)}\leq C\Big{(}\tau\sum_{k=1}^{l}\|d_{\tau}u^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}+\tau\sum_{k=1}^{l}\|\partial_{x}u^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}\Big{)}$ (3.48)
and the result follows from Lemma 3.5. . ∎
3.9. Lemma (energy estimates for the concentration)
Let assumptions (3.3) hold. Then
$\max_{l=0,...,K}\|c^{l}\|^{2}_{\operatorname{L}_{2}(\Omega)}+\tau\sum_{k=1}^{K}\|c^{k}\|^{2}_{\operatorname{H}^{1}(\Omega)}+\tau\sum_{k=1}^{K}\|d_{\tau}c^{k}\|^{2}_{(\operatorname{H}^{1}(\Omega))^{\prime}}\leq C.$ (3.49)
Chứng minh.
Setting $\chi=c^{k}$ in (3.1c) gives
$(d_{\tau}c^{k},c^{k})+(D(u^{k})\nabla c^{k},\nabla c^{k})=-(d_{\tau}u^{k},c^{k})-(D_{1}(u^{k},c^{k-1})\nabla\phi^{k},\nabla c^{k})$ (3.50)
and noting (3.26), (3.9) and (3.11), we obtain
$\frac{\tau}{2}\|d_{\tau}c^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}+\frac{1}{2}d_{\tau}\|c^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}\!\!\!$ $+$ $\!\!\!D_{\min}|c^{k}|^{2}_{\operatorname{H}^{1}(\Omega)}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \frac{1}{2}\|d_{\tau}u^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}+\frac{1}{2}\|c^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}$ (3.51)
$+C|\phi^{k}|^{2}_{\operatorname{H}^{1}(\Omega)}+\frac{D_{\min}}{2}|c^{k}|^{2}_{\operatorname{H}^{1}(\Omega)}.$
After rearranging (3.51), multiplying the resulting inequality by $\tau$ and summing over $k=1,2,...,l$, for any $l\leq K$, we have
$\frac{\tau^{2}}{2}\sum_{k=1}^{l}\|d_{\tau}c^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}+\frac{1}{2}\|c^{l}\|^{2}_{\operatorname{L}_{2}(\Omega)}+\tau\frac{D_{\min}}{2}\sum_{k=1}^{l}|c^{k}|^{2}_{\operatorname{H}^{1}(\Omega)}\\ \leq\frac{1}{2}\|c^{0}\|^{2}_{\operatorname{L}_{2}(\Omega)}+\frac{\tau}{2}\sum_{k=1}^{l}\|d_{\tau}u^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}+\frac{\tau}{2}\sum_{k=1}^{l}\|c^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}+\tau C\sum_{k=1}^{l}|\phi^{k}|^{2}_{\operatorname{H}^{1}(\Omega)}.$
Thus, noting from assumption (3.3) that $c_{0}\in\operatorname{L}_{2}(\Omega)$, we have
12($\frac{1}{2}\left(1-\tau\right)\|c^{l}\|^{2}_{\operatorname{L}_{2}(\Omega)}\!\!\!$ $+$ $\!\!\!\tau\frac{D_{\min}}{2}\sum_{k=1}^{l}|c^{k}|^{2}_{\operatorname{H}^{1}(\Omega)}\leq C\Big{(}1+\tau\sum_{k=1}^{l}\|d_{\tau}u^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}$ (3.52)
$+\leavevmode\nobreak\ \tau\sum_{k=1}^{l}|\phi^{k}|^{2}_{\operatorname{H}^{1}(\Omega)}+\tau\sum_{k=1}^{l-1}\|c^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}\Big{)}.$
Since (3.3) holds we have $\tau<1/2$ and hence using the discrete Gronwall Lemma, we obtain
$\begin{split}\frac{1}{4}\max_{l=1,...,K}\|c^{l}\|^{2}_{\operatorname{L}_{2}(\Omega)}+\tau\frac{D_{\min}}{2}\ \sum_{k=1}^{K}|c^{k}|^{2}_{\operatorname{H}^{1}(\Omega)}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \\ \leq C\Big{(}1+\tau\sum_{k=1}^{K}\|d_{\tau}u^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}+\tau\sum_{k=1}^{K}|\phi^{k}|^{2}_{\operatorname{H}^{1}(\Omega)}\Big{)}\operatorname{e}\!^{\frac{T}{2}}\end{split}$ (3.53)
and from lemmas 3.5 and 3.8 we conclude
$\frac{1}{4}\max_{l=0,...,K}\|c^{l}\|^{2}_{\operatorname{L}_{2}(\Omega)}+\tau\sum_{k=1}^{K}|c^{k}|^{2}_{\operatorname{H}^{1}(\Omega)}\leq C.$ (3.54)
To prove the last estimate in (3.49), we note from (3.1c), (3.9) and (3.11) that
$(d_{\tau}c^{k},\chi)_{((\operatorname{H}^{1}(\Omega))^{\prime},\operatorname{H}^{1}(\Omega))}\leq\|D(u^{k})\nabla c^{k}\|_{\operatorname{L}_{2}(\Omega)}\|\nabla\chi\|_{\operatorname{L}_{2}(\Omega)}+\|d_{\tau}u^{k}\|_{\operatorname{L}_{2}(\Omega)}\|\chi\|_{\operatorname{L}_{2}(\Omega)}\\ +\|D_{1}(u^{k},c^{k-1})\nabla\phi^{k}\|_{\operatorname{L}_{2}(\Omega)}|\nabla\chi\|_{\operatorname{L}_{2}(\Omega)}\\ \leq C(|c^{k}|_{\operatorname{H}^{1}(\Omega)}+|\phi^{k}|_{\operatorname{H}^{1}(\Omega)}+\|d_{\tau}u^{k}\|_{\operatorname{L}_{2}(\Omega)})\|\chi\|_{\operatorname{H}^{1}(\Omega)}$ (3.55)
and so we conclude
$\|d_{\tau}c^{k}\|_{(\operatorname{H}^{1}(\Omega))^{\prime}}^{2}\leq C(|c^{k}|^{2}_{\operatorname{H}^{1}(\Omega)}+|\phi^{k}|^{2}_{\operatorname{H}^{1}(\Omega)}+\|d_{\tau}u^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}).$ (3.56)
The required result follows after multiplying by $\tau$, summing over $k=1,2,...,K$ and noting (3.54) and lemmas 3.5 and 3.8. ∎
Proof  of Theorem 3.1 Theorem 3.1 is a direct consequence of lemmas 3.4, 3.6 and 3.7.   ∎
4. Weak convergence of the limits
In this section, we prove that as $\tau\to 0$ subsequences of the solution to (3.1)–(3.1c) converge to a solution of (2.11a)–(2.11c).
Introduce the notation
$u^{-}_{\tau}(t)=u^{k-1},\quad u^{+}_{\tau}(t)=u^{k}\text{ for }t_{k-1}<t\leq t_{k},$ (4.1)
and do the same $c^{\pm}_{\tau}(t),\hat{c}_{\tau}(t),\phi^{\pm}_{\tau}(t),\hat{\phi}_{\tau}(t)$. Also use $u_{\tau}$ to denote the interpolate
$\hat{u}_{\tau}(t)=\frac{t-t_{k-1}}{\tau}u^{k}+\frac{t_{k}-t}{\tau}u^{k-1},\text{ for }t_{k-1}\leq t\leq t_{k},k={0},\dotsc,{K}.$ (4.2)
Next recalling (3.2), the terms on the left hand side of (3.1) can be reformulated as
$\begin{split}((m(c^{k-1}))^{-}h_{1}^{\prime}(u^{k}),\chi)+((m(c^{k-1}))^{+}h_{2}^{\prime}(u^{k}),\chi)=30\left(m(c^{k-1})u^{k-1}u^{k}(u^{k}-1)^{2},\chi\right)\\ +30\left((m(c^{k-1}))^{+}(u^{k}-1)u^{k}(u^{k-1}-u^{k}),\chi\right).\end{split}$ (4.3)
Thus rewriting (3.1)–(3.1b) and using the above notation, gives, for $t^{k-1}<t\leq t^{k}$, $k=1,\ldots,K$, and for all $(\chi,\eta)\in\operatorname{H}^{1}(\Omega)\times\operatorname{H}^{1}_{0|\Gamma_{1}\cup\Gamma_{2}}(\Omega)$
$(\partial_{t}\hat{u}_{\tau},\chi)+(\boldsymbol{{A}}(\nabla u^{+}_{\tau}(t))\nabla u^{+}_{\tau}(t),\nabla\chi)+(g^{\prime}(u^{+}_{\tau}(t)),\chi)\\ =({\mathcal{F}}(c_{\tau}^{-}(t),u_{\tau}^{-}(t),u_{\tau}^{+}(t)),\chi)$ (4.4a)
$(\partial_{t}\hat{c}_{\tau},\chi)+(D(u^{+}_{\tau}(t))\nabla c^{+}_{\tau}(t),\nabla\chi)\\ =-(\hat{u}_{\tau}^{\prime}(t),\chi)-(D_{1}(u^{+}_{\tau}(t),c^{-}_{\tau}(t))\nabla\phi^{+}_{\tau}(t),\nabla\chi)$ (4.4b)
and
$(\sigma(u^{+}_{\tau}(t))\nabla\phi^{+}_{\tau}(t),\nabla\eta)=-(\hat{u}_{\tau}^{\prime}(t)+\sigma^{\prime}(u_{\tau}^{+}(t))\partial_{x}u_{\tau}^{+}(t),\eta),$ (4.4c)
where
$\begin{split}({\mathcal{F}}(c_{\tau}^{-}(t),u_{\tau}^{-}(t),u_{\tau}^{+}(t)),\chi):=30\left(m(c_{\tau}^{-})u_{\tau}^{-}(t)u_{\tau}^{+}(t)(u_{\tau}^{+}(t)-1)^{2},\chi\right)\\ +30\left((m(c_{\tau}^{-}(t)))^{+}(u_{\tau}^{+}(t)-1)u_{\tau}^{+}(t)(u_{\tau}^{-}(t)-u_{\tau}^{+}(t)),\chi\right).\end{split}$
4.1. Lemma (stability bounds)
Assuming (3.3), we have
$\|u_{\tau}^{+}\|_{\operatorname{L}_{\infty}([0,T];\operatorname{L}_{\infty}(\Omega))}+\|u_{\tau}^{+}\|_{\operatorname{L}_{2}([0,T];\operatorname{H}^{1}(\Omega))}+\|\partial_{t}\hat{u}_{\tau}\|_{\operatorname{L}_{2}([0,T];\operatorname{L}_{2}(\Omega))}\leq C,$ (4.5)
$\|c_{\tau}^{+}\|_{\operatorname{L}_{\infty}([0,T];\operatorname{L}_{2}(\Omega))}+\|c_{\tau}^{+}\|_{\operatorname{L}_{2}([0,T];\operatorname{H}^{1}(\Omega))}+\|\partial_{t}\hat{c}_{\tau}\|_{\operatorname{L}_{2}([0,T];(\operatorname{H}^{1}(\Omega))^{\prime})}\leq C,$ (4.6)
$\|\phi_{\tau}^{+}\|_{\operatorname{L}_{2}([0,T];\operatorname{H}^{1}(\Omega))}\leq C.$ (4.7)
Chứng minh.
The first and second bounds in (4.5) follow from lemmas 3.2 and 3.5. For the last bound in (4.5), we note from (4.2) that $\partial_{t}\hat{u}_{\tau}(t)=d_{\tau}u^{k}$ on $(t_{k-1},t_{k})$, $k=1,2,...,K$ and thus the result follows from Lemma 3.5.
The bounds in (4.6) and (4.7) follow using similar arguments and the results of lemmas 3.9 and 3.8 respectively. ∎
4.2. Lemma (weak compactness)
There exist $(u,c,\phi)$ and $(\partial_{t}u,\partial_{t}c)$ such that for a sequence $(\tau_{n})_{n\in\mathbb{N}}$ of positive numbers with $\tau_{n}\to 0$ as $n\to\infty$, we have the following for $1\leq p<\infty$ if dim $\text{dim }\Omega=2$ and $61\leq p<$ 6 if dim $\text{dim }\Omega=3$, up to a subsequence, which for simplicity of presentation we again denote by $\tau_{n}$,
$\hat{u}_{\tau_{n}},u_{\tau_{n}}^{\pm}\overset{\ast}{\rightharpoonup}u\text{ in }\operatorname{L}_{\infty}([0,T];\operatorname{L}_{\infty}(\Omega)),$ (4.8a)
$\hat{u}_{\tau_{n}},u_{\tau_{n}}^{\pm}\rightharpoonup u\text{ in }\operatorname{L}_{2}([0,T];\operatorname{H}^{1}(\Omega)),$ (4.8b)
$\partial_{t}\hat{u}_{\tau_{n}}\rightharpoonup\partial_{t}u\text{ in }\operatorname{L}_{2}([0,T];\operatorname{L}_{2}(\Omega)),$ (4.8c)
$\hat{c}_{\tau_{n}},c_{\tau_{n}}^{\pm}\overset{\ast}{\rightharpoonup}c\text{ in }\operatorname{L}_{\infty}([0,T];\operatorname{L}_{2}(\Omega)),$ (4.8d)
$\hat{c}_{\tau_{n}},c_{\tau_{n}}^{\pm}\rightharpoonup c\text{ in }\operatorname{L}_{2}([0,T];\operatorname{H}^{1}(\Omega)),$ (4.8e)
$\partial_{t}\hat{c}_{\tau_{n}}\rightharpoonup\partial_{t}c\text{ in }\operatorname{L}_{2}([0,T];(\operatorname{H}^{1}(\Omega))^{\prime}),$ (4.8f)
$\phi_{\tau_{n}}^{\pm}\rightharpoonup\phi\text{ in }\operatorname{L}_{2}([0,T];\operatorname{H}^{1}(\Omega)),$ (4.8g)
$u_{\tau_{n}}^{\pm}\rightarrow u\text{ in }\operatorname{L}_{2}([0,T];\operatorname{L}_{p}(\Omega)),$ (4.8h)
$c_{\tau_{n}}^{\pm}\rightarrow c\text{ in }\operatorname{L}_{2}([0,T];\operatorname{L}_{p}(\Omega)).$ (4.8i)
Chứng minh.
From Lemma 4.1 it follows that there are weakly convergent subsequences that converge to the appropriate limits in (4.8a)-(4.8g). To establish (4.8h) and (4.8i) we use (4.8b), (4.8c), (4.8e), (4.8f) and the Aubin-Lions Lemma (e.g. Roubiček, 2005). ∎
4.3. Lemma (strong convergence)
For $u\in\operatorname{L}_{\infty}([0,T];\operatorname{L}_{\infty}(\Omega))$ and $c\in\operatorname{L}_{2}([0,T];\operatorname{L}_{q}(\Omega))$, with $1\leq q<\infty$ if dim $\text{dim }\Omega=2$ and $31\leq q<$ 3 if dim $\text{dim }\Omega=3$, we have the following, for $1\leq p<\infty$ if dim $\text{dim }\Omega=2$ and $61\leq p<$ 6 if dim $\text{dim }\Omega=3$, as $\tau_{n}\to 0$
$g^{\prime}(u_{\tau_{n}}^{+})\to g^{\prime}(u)\text{ in }\operatorname{L}_{2}([0,T];\operatorname{L}_{p}(\Omega)),$ (4.9a)
$u_{\tau}^{-}u_{\tau}^{+}(u_{\tau}^{+}-1)^{2}\to u^{2}(u-1)^{2}\text{ in }\operatorname{L}_{2}([0,T];\operatorname{L}_{p}(\Omega)),$ (4.9b)
$D(u_{\tau_{n}}^{+})\rightarrow D(u)\text{ in }\operatorname{L}_{2}([0,T];\operatorname{L}_{p}(\Omega)),$ (4.9c)
$\sigma(u_{\tau_{n}}^{+})\rightarrow\sigma(u)\text{ in }\operatorname{L}_{2}([0,T];\operatorname{L}_{p}(\Omega)),$ (4.9d)
$m(c^{-}_{\tau_{n}})\rightarrow m(c)\text{ in }\operatorname{L}_{2}([0,T];\operatorname{L}_{p}(\Omega)),$ (4.9e)
$D_{1}(u^{+}_{\tau_{n}},c^{-}_{\tau_{n}})\to D_{1}(u,c)\text{ in }\operatorname{L}_{2}([0,T];\operatorname{L}_{q}(\Omega)),$ (4.9f)
$30m(c_{\tau}^{-})u_{\tau}^{-}u_{\tau}^{+}(u_{\tau}^{+}-1)^{2}\to m(c)h^{\prime}(u)\text{ in }\operatorname{L}_{2}([0,T];\operatorname{L}_{q}(\Omega)),$ (4.9g)
$(m(c_{\tau}^{-}))^{+}(u_{\tau}^{+}-1)u_{\tau}^{+}(u_{\tau}^{-}-u_{\tau}^{+})\to 0\text{ in }\operatorname{L}_{2}([0,T];\operatorname{L}_{2}(\Omega)).$ (4.9h)
Chứng minh.
Noting that $g^{\prime}(u),D(u),\sigma(u)$ are polynomials in $u$ and $m(c)$ is linear in $c$, the proofs of (4.9a) – (4.9e) follow from Lemma 4.1, (4.8a), (4.8d), (4.8h) and (4.8h). By the triangle inequality, the generalized Hölder inequality with $1/q=1/p+1/p^{\prime}$, (4.6) and (3.9) we obtain
$\|D(u_{\tau}^{+})c_{\tau}^{-}-D(u)c\|_{\operatorname{L}_{2}([0,T];\operatorname{L}_{q}(\Omega))}\\ =\|(D(u_{\tau}^{+})-D(u))c_{\tau}^{-}+D(u)(c_{\tau}^{-}-c)\|_{\operatorname{L}_{2}([0,T];\operatorname{L}_{q}(\Omega))}\\ \leq\|c_{\tau}^{-}\|_{\operatorname{L}_{2}([0,T];\operatorname{L}_{p^{\prime}}(\Omega))}\|D(u_{\tau}^{+})-D(u)\|_{\operatorname{L}_{2}([0,T];\operatorname{L}_{p}(\Omega))}\\ +\|D(u)\|_{\operatorname{L}_{2}([0,T];\operatorname{L}_{p^{\prime}}(\Omega))}\|c_{\tau}^{-}-c\|_{\operatorname{L}_{2}([0,T];\operatorname{L}_{p}(\Omega))}\\ \leq C(\|D(u_{\tau}^{+})-D(u)\|_{\operatorname{L}_{2}([0,T];\operatorname{L}_{p}(\Omega))}+\|c_{\tau}^{-}-c\|_{\operatorname{L}_{2}([0,T];\operatorname{L}_{p}(\Omega))})$ (4.10)
and (4.9f) then follows by noting the definition of $D_{1}(u,c)$ in (2.10), and using (4.9c) and (4.8i). Adopting similar techniques and noting from (2.12) that $h^{\prime}(u)=30u^{2}(u-1)^{2}$, from (3.11), (4.5), (4.9b) and (4.9e) we conclude (4.9g).
Noting from (4.2) that $\tau\partial_{t}\hat{u}_{\tau}=(u_{\tau}^{-}-u_{\tau}^{+})$ for $t\in(t^{k-1},t^{k})$, from (3.11) and (4.5) we have
$\|(m(c_{\tau}^{-}))^{+}(u_{\tau}^{+}-1)u_{\tau}^{+}(u_{\tau}^{-}-u_{\tau}^{+})\|_{\operatorname{L}_{2}([0,T];\operatorname{L}_{2}(\Omega))}\\ \leq C\|u_{\tau}^{+}-u_{\tau}^{-}\|_{\operatorname{L}_{2}([0,T];\operatorname{L}_{2}(\Omega))}=C\tau$ (4.11)
which concludes the proof. ∎
4.4. Lemma (weak convergence)
For $D(u)\nabla c$, $D_{1}(u,c)\nabla\phi$, $\sigma(u)\nabla\phi$, $\boldsymbol{{A}}(\nabla u)\nabla u\in\operatorname{L}_{2}([0,T];(\operatorname{L}_{2}(\Omega))^{d})$ we have the following as $\tau_{n}\to 0$
$D(u^{+}_{\tau_{n}})\nabla c^{+}_{\tau_{n}}\rightharpoonup D(u)\nabla c\text{ in }\operatorname{L}_{2}([0,T];(\operatorname{L}_{2}(\Omega))^{d}),$ (4.12a)
$D_{1}(u^{+}_{\tau_{n}},c^{-}_{\tau_{n}})\nabla\phi^{+}_{\tau_{n}}\rightharpoonup D_{1}(u,c)\nabla\phi\text{ in }\operatorname{L}_{2}([0,T];(\operatorname{L}_{2}(\Omega))^{d}),$ (4.12b)
$\sigma(u^{+}_{\tau_{n}})\nabla\phi_{\tau_{n}}^{+}\rightharpoonup\sigma(u)\nabla\phi\text{ in }\operatorname{L}_{2}([0,T];(\operatorname{L}_{2}(\Omega))^{d}),$ (4.12c)
$\boldsymbol{{A}}(\nabla u^{+}_{\tau_{n}})\nabla u^{+}_{\tau_{n}}\rightharpoonup\boldsymbol{{A}}(\nabla u)\nabla u\text{ in }\operatorname{L}_{2}([0,T];(\operatorname{L}_{2}(\Omega))^{d}).$ (4.12d)
Chứng minh.
From Lemma 4.1 we have
$\int_{0}^{T}\int_{\Omega}D(u^{+}_{\tau_{n}})\nabla c^{+}_{\tau_{n}}\cdot\nabla c^{+}_{\tau_{n}}=\int_{0}^{T}\int_{\Omega}|D^{1/2}(u^{+}_{\tau_{n}})\nabla c^{+}_{\tau_{n}}|^{2}\\ \leq\|D^{1/2}(u^{+}_{\tau_{n}})\nabla c^{+}_{\tau_{n}}\|_{\operatorname{L}_{2}([0,T];(\operatorname{L}_{2}(\Omega))^{d})}^{2}\leq D_{\max}\|c^{+}_{\tau_{n}}\|_{\operatorname{L}_{2}([0,T];\operatorname{H}^{1}(\Omega))}^{2}\leq C$ (4.13)
and hence it follows that there exists a $\xi_{1}\in\operatorname{L}_{2}([0,T];(\operatorname{L}_{2}(\Omega))^{d})$ such that
$D(u^{+}_{\tau_{n}})\nabla c^{+}_{\tau_{n}}\rightharpoonup\xi_{1}\text{ in }\operatorname{L}_{2}([0,T];(\operatorname{L}_{2}(\Omega))^{d}).$ (4.14)
We now show that $\xi_{1}=D(u)\nabla c$, to this end we note that
$\Big{|}\int_{0}^{T}\int_{\Omega}\Big{(}D(u_{\tau_{n}}^{+})\nabla c_{\tau_{n}}^{+}-D(u)\nabla c\Big{)}\cdot\nabla\psi\Big{|}\\ =\Big{|}\int_{0}^{T}\int_{\Omega}\Big{(}D(u_{\tau_{n}}^{+})\nabla c_{\tau_{n}}^{+}-D(u)\nabla c_{\tau_{n}}^{+}+D(u)\nabla c_{\tau_{n}}^{+}-D(u)\nabla c\Big{)}\cdot\nabla\psi\Big{|}\\ \leq\Big{|}\int_{0}^{T}\int_{\Omega}D(u)\Big{(}\nabla c_{\tau_{n}}^{+}-\nabla c\Big{)}\cdot\nabla\psi\Big{|}+\Big{|}\int_{0}^{T}\int_{\Omega}\Big{(}D(u^{+}_{\tau_{n}})-D(u)\Big{)}\nabla c_{\tau_{n}}^{+}\cdot\nabla\psi\Big{|}\\ \leq\Big{|}\int_{0}^{T}\int_{\Omega}D(u)\Big{(}\nabla c_{\tau_{n}}^{+}-\nabla c\Big{)}\cdot\nabla\psi\Big{|}\\ +\|D(u^{+}_{\tau_{n}})-D(u)\|_{\operatorname{L}_{2}([0,T];\operatorname{L}_{2}(\Omega))}\|\nabla c_{\tau_{n}}^{+}\cdot\nabla\psi\|_{\operatorname{L}_{2}([0,T];\operatorname{L}_{2}(\Omega))}.$ (4.15)
Since $\|D(u)\|_{\operatorname{L}_{\infty}(\Omega)}\leq C$ from (4.8e), (4.9c) and (4.6) we conclude (4.12a). Using similar techniques we can obtain (4.12b) and (4.12c). We omit the proof of (4.12d) as it is given in (Burman and Rappaz, 2003, Lemma 5.4, p1152). ∎
4.5. Proof of Theorem 2.8
Now we can prove our main result, Theorem 2.8.
Chứng minh.
In Theorem 3.1 we established the existence of a solution $(u^{k},c^{k},\phi^{k})\in(\operatorname{H}^{1}(\Omega))^{2}\times\operatorname{H}^{1}_{0|\Gamma_{1}\cup\Gamma_{2}}(\Omega)$ of (3.1)–(3.1c). From lemmas 4.24.4 we conclude the existence of a solution $(u,c,\phi)\in(\operatorname{H}^{1}(\Omega))^{2}\times\operatorname{H}^{1}_{0|\Gamma_{1}\cup\Gamma_{2}}(\Omega)$ of (2.11a)–(2.11c), with $(u(0),c(0))=(u_{0},c_{0})\in\operatorname{H}^{1}(\Omega)\times\operatorname{L}_{2}(\Omega)$ and $u\in\operatorname{L}_{2}([0,T];\operatorname{H}^{1}(\Omega))\cap\operatorname{H}^{1}([0,T];\operatorname{L}_{2}(\Omega))\cap\operatorname{L}_{\infty}([0,T];\operatorname{L}_{\infty}(\Omega))$, $c\in\operatorname{L}_{2}([0,T];\operatorname{H}^{1}(\Omega))\cap\operatorname{H}^{1}([0,T];(\operatorname{H}^{1}(\Omega))^{\prime})\cap\operatorname{L}_{\infty}([0,T];\operatorname{L}_{2}(\Omega))$ and $\phi\in\operatorname{L}_{2}([0,T];\operatorname{H}^{1}_{0|\Gamma_{1}\cup\Gamma_{2}}(\Omega))$. ∎
5. Numerical Simulations
In this section we present a finite element discretization of (2.11a)-(2.11c) together with numerical examples that illustrate how dendritic structures can arise from the model.
5.1. Finite element approximation
Until now, for ease of presentation, we have considered a simplified version of the model from Chen et al. (2015) in which, where possible, all physical parameters were set to $1$. However in this section, in order to produce computational simulations in which dendritic patterns arise, we consider the original model from Chen et al. (2015) that has the physical parameters reinserted.
As in Section 3, we partition the time interval $[0,T]$ into $K$ equidistant steps, $0=t_{0}<t_{1}<\ldots<t_{K-1}<t_{K}=T$, so that for $k=1,\ldots,K$, $t_{k}:=k\tau$. We let $\mathscr{T}_{h}$ be a regular triangulation of $\Omega$ into disjoint open simplices $\mathcal{T}$ and we define $h:=\max_{\mathcal{T}\in\mathscr{T}_{h}}\operatorname{diam}{\mathcal{T}}$ the maximal element size of $\mathscr{T}_{h}$.
Associated with ${\mathscr{T}}_{h}$ are the finite element spaces
$V_{h}=\{v\in\operatorname{C}^{0}(\bar{\Omega});\leavevmode\nobreak\ v|_{\mathcal{T}}\in P_{2}(\mathcal{T}),\leavevmode\nobreak\ \forall\leavevmode\nobreak\ \mathcal{T}\in\mathscr{T}_{h}\}\subset\operatorname{H}^{1}(\Omega),$
$V_{0,h}=\{v\in\operatorname{C}^{0}(\bar{\Omega});\leavevmode\nobreak\ v=0\text{ on }\Sigma\subseteq\partial\Omega;\leavevmode\nobreak\ v|_{\mathcal{T}}\in P_{2}(\mathcal{T}),\leavevmode\nobreak\ \forall\leavevmode\nobreak\ \mathcal{T}\in\mathscr{T}_{h}\}\subset\operatorname{H}^{1}_{0|\Sigma}(\Omega),$
where $\Sigma=\Gamma_{1}\cup\Gamma_{2}$ and $P_{2}(\mathcal{T})$ denotes the set of polynomials of maximum degree $2$ on $\mathcal{T}$.
Using a standard Galerkin finite element method in space and a semi-implicit Euler discretization in time we obtain the following discretisation of a variant of (2.11a)-(2.11c) in which the physical parameters $\epsilon,\gamma,\nu,\mu$ from Chen et al. (2015) have been reinstated, the boundedness imposed on the $D_{1}(u,c)$, see (2.10), is removed and the electrochemical reaction rate $m(c)$ is replaced with a modified version $\tilde{m}(c)$ to highlight its exponential structure, see (Chen et al., 2015, Appendix B):
Set $u^{0}_{h}=u_{0}(\boldsymbol{x})$ and $c^{0}_{h}=c_{0}(\boldsymbol{x})$, then for $k=1,2\ldots,K$, given $(u^{k-1}_{h},c^{k-1}_{h})\in[V_{h}]^{2}$ find $(u^{k}_{h},c^{k}_{h},\phi^{k}_{h})\in[V_{h}]^{2}\times V_{0,h}$ such that for all $(\chi_{h},\eta_{h})\in V_{h}\times V_{0,h}$
$\epsilon^{2}(d_{\tau}u^{k}_{h},\chi_{h})+(\boldsymbol{{A}}(\nabla u^{k-1}_{h})\nabla u^{k}_{h},\nabla\chi_{h})+\gamma(g^{\prime}(u^{k-1}_{h}),\chi_{h})=(\tilde{m}(c^{k-1}_{h})h^{\prime}(u^{k-1}_{h}),\chi_{h}),$ (5.1a)
$(\sigma(u^{k}_{h})\nabla\phi^{k}_{h},\nabla\eta_{h})=-\nu(d_{\tau}u^{k}_{h},\eta_{h}),$ (5.1b)
$(d_{\tau}c^{k}_{h},\chi_{h})+(D(u^{k}_{h})\nabla c^{k},\nabla\chi_{h})+(D(u^{k}_{h})c^{k-1}_{h}\nabla\phi^{k}_{h},\nabla\chi_{h})=-\mu(d_{\tau}u^{k}_{h},\chi_{h}),$ (5.1c)
where $\tilde{m}(c)=C_{1}(e^{-C_{2}}-ce^{C_{2}})$. The physical parameters are defined, together with their values in the computations, in Table 1. We note that (5.1a) – (5.1c) are linear in $u^{k}$, $\phi^{k}$ and $c^{k}$ respectively.
The computational results we present are produced by implementing a standard adaptive strategy, in particular, we use a gradient-based indicator which involves the gradient of the solution of the phase-field equation and the reaction-diffusion equation, given by
$\eta_{\mathcal{T}}:=\sqrt{{\left|\nabla u^{k}_{h}\right|}^{2}_{\mathcal{T}}+{\left|\nabla c^{k}_{h}\right|}^{2}_{\mathcal{T}}},$ (5.2)
such that simplicies on which $\eta_{\mathcal{T}}\geq 160$ are refined and simplices on which $\eta_{\mathcal{T}}\leq 80$ are coarsened.
5.2. Numerical examples
In this section we present numerical simulations of dendritic crystal growth.
[…]
$u_{0}(x_{1},x_{2})=\Bigg{\{}\begin{array}[]{ll}1,&(x_{1}-r_{0}-\sqrt{r_{0}^{2}-x_{2}^{2}})(x_{1}+r_{0}+\sqrt{r_{0}^{2}-x_{2}^{2}})(x_{2}^{2}-r_{0}^{2})\geq 0,\\ 0,&\text{elsewhere}.\end{array}$ (5.3)
and
$c_{0}(x_{1},x_{2})=\Bigg{\{}\begin{array}[]{ll}0,&x_{1}=0,\\ 1,&\text{elsewhere.}\end{array}$ (5.4)
In the simulations presented below we set $r_{0}=0.1$.
For the implementation of the numerical scheme we used the DUNE Python module, Dedner and Nolte (2018), and the DUNE Alugrid module, Alkämper et al. (2015). Since the solution is symmetric in the $x_{2}$-direction, for the sake of computational cost, we reduced our computations from the rectangular domain $\Omega=[0,9]\times[-9,9]$ to the square domain $\Omega^{\prime}=[0,9]^{2}$. The values of the parameters used in the simulations are presented in Table 1.
We consider the polar form of the anisotropy function (2.15):
$\tilde{a}(\theta)=a_{0}(1+4\delta\cos(4\theta)),$ (5.5)
with $\theta=\arctan(\partial_{x_{2}}u/\partial_{x_{1}}u)$, which yields the anisotropy tensor
$\boldsymbol{{A}}(\theta)=\begin{bmatrix}\tilde{a}^{2}(\theta)&-\tilde{a}(\theta)\tilde{a}^{\prime}(\theta)\\ \tilde{a}(\theta)\tilde{a}^{\prime}(\theta)&\tilde{a}^{2}(\theta)\end{bmatrix}.$ (5.6)
We present four sets of results, Figure 1 – Figure 4. In Figure 1 we show a simulation of the evolution of a dendritic structure, while in Figures 2 and 3 we examine how the dendritic structure is affected by variations in the anisotropy strength $\delta$ and in the parameter $\mu$ respectively. In Figure 4 we investigate the effect adding stochastic noise has on the dendritic structure.
Figure 1 depicts the numerical approximations of the order parameter, $u_{h}^{k}$ (top row), the concentration of lithium atoms, $c_{h}^{k}$ (middle row), and the electrostatic potential, $\phi_{h}^{k}$ (bottom row), each displayed at $t_{k}=0.061$ (left) $t_{k}=0.244$ (middle) and $t_{k}=0.427$ (right). In these results we set $\delta=0.05$ and $\mu=2$. From the results for the order parameter we clearly see the formation of a dendritic structure, while a more diffuse form of this structure is mirrored in the concentration results and a very diffuse underlying dendritic structure can also be observed in the electric potential results, with the region, in which ${\left|\phi^{k}_{h}\right|}$ attains its maximum value, growing as the dendrite grows, thus eventually causing the battery to fail. When comparing the dendritic structure that forms in Figure 1 to the dendritic structures that are displayed in Figure 4 in Chen et al. (2015) and Figure 3a in Mu et al. (2019) we see that in Figure 1 the structure is smooth and does not form the side branches perpendicular to the growth direction of the dendritic structure, that can be observed in Chen et al. (2015) and Mu et al. (2019). In Figures 2 and 4 below, we see that dendritic structures with side branches can be observed by increasing the anisotropy strength, $\delta$, or adding stochastic noise.
In Figure 2 we examine how the anisotropy strength, $\delta$, affects the formation of dendritic crystals. To this end we recall the bound $\delta<1/15\approx 0.067$ that we impose in assumptions (3.3), recall (3.3), in order for our theoretical results to hold. The specific nature of the bound can be understood from the proof of Lemma 2.5. Figure 2 displays plots of the numerical approximation of the order parameter, $u_{h}^{k}$, at $t_{k}=0.366$ for $\mu=2$ and six values of $\delta$. From this figure we see that the anisotropy strength has a notable impact on the rate at which the dendritic structure grows, in particular the speed of growth increases as $\delta$ decreases. In addition, we see that the value $\delta=1/15\approx 0.067$ plays a significant role; in the simulations where $\delta\ll 1/15\approx 0.067$, i.e. Figure 2(a) – Figure 2(d) in which $\delta=0,0.01,0.02,0.05$ respectively, the geometry of the dendritic structure is smooth, while in Figure 2(e), where $\delta=0.065$ and hence the condition $\delta<1/15$ is only just satisfied, the structure is no longer smooth and small instabilities, perpendicular to the growth of the dendritic structure, have started to form. The instabilities are more pronounced in Figure 2(f), which shows $\delta=0.1$ which is significantly greater than $\delta=1/15$. $\delta\geq 1/15$, the numerical scheme still produces solutions for values of $\delta>1/15$.
In Figure 3 we examine how the evolution of the order parameter, $u$, varies with the parameter $\mu$, which is related to the ratio of site density of Li and the standard bulk concentration of electrolyte solution as described in Chen et al. (2015). We display plots of the numerical approximation of the order parameter, $u_{h}^{k}$, at $t_{k}=0.122$ for $\delta=0.05$ and five values of $\mu$, together with, for comparison, a plot of $u_{h}^{k}$ resulting from solving the system (5.1a)–(5.1c) with $\delta=0$ and $\mu=0$. From this figure we see that for small values of $\mu$, i.e. $\mu=0$ and $\mu=0.5$, no dendritic structures arise and instead the geometry takes a form that is relatively radially symmetric, while for larger values of $\mu$, namely $\mu=1,1.5,2$ dendritic patterns arise, with the dendritic structure becoming more pronounced, yet growing more slowly, as $\mu$ increases.
As observed above, in Figure 1 no side branches are formed perpendicular to the growth direction of the dendritic structure, whereas in the results in Chen et al. (2015) and Mu et al. (2019) side branches do form. From Figure 2 we see that side branches are observed if the anisotropy strength, $\delta$, is taken large enough. An alternative approach to yield the formation of side branches is to add stochastic noise to the model, to this end we follow the authors in Kobayashi (1993) and add a stochastic term, $\beta u_{h}^{k}(1-u^{k}_{h})X$, to (5.1a), so that it becomes
$\epsilon^{2}(d_{\tau}u^{k}_{h},\chi_{h})+(\boldsymbol{{A}}(\nabla u^{k-1}_{h})\nabla u^{k}_{h},\nabla\chi_{h})+\gamma(g^{\prime}(u^{k-1}_{h}),\chi_{h})=(\tilde{m}(c^{k-1}_{h})h^{\prime}(u^{k-1}_{h}),\chi_{h})$
$\hskip 85.35826pt-(\beta u_{h}^{k}(1-u^{k}_{h})X,\chi_{h}),\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\leavevmode\nobreak\ \chi_{h}\in V_{h},$ (5.7)
where $X\sim\mathcal{U}(-1/2,1/2)$ and $\beta>0$ is a scalar that controls the noise strength. Results obtained from the model with (5.1a) replaced by (5.7), and $\delta=0.05$ and $\mu=2$, are shown in Figure 4, which displays plots of the numerical approximation of the order parameter, $u_{h}^{k}$, at $t_{k}=0.305$ for four values of $\beta$. From this figure we see that when $\beta=0.03$ side branches have started to form and that the length of these branches increases as $\beta$ increases.
Bảng 1. Dimensionless parameters.
Parameter Notation Value
Diffusion coefficient in the electrolyte $D^{\text{e}}$ $3.0$
Diffusion coefficient in the electrode $D^{\text{s}}$ $2.5$
Conductivity coefficient in the electrolyte $\sigma^{\text{e}}$ $100.0$
Conductivity coefficient in the electrode $\sigma^{\text{s}}$ $10^{7}$
Interfacial mobility $\gamma$ $1.0$
Characteristic time scale $\epsilon^{2}$ $0.0003$
Anisotropy strength $\delta$ $0.05$ (unless otherwise specified)
Interfacial thickness $a_{0}$ $0.01$
Electrochemical reaction rate coefficients $C_{1}$ $C_{2}$ $1/15$ $1.9345$
Coefficient multiplying $d_{\tau}u_{h}^{k}$ in (5.1b) $\nu$ $76.4$
Coefficient multiplying $d_{\tau}u_{h}^{k}$ in (5.1c) $\mu$ $2.0$ (unless otherwise specified)
Hình 1. Dendritic formation with $\delta=0.05$ and $\mu=2$: $u_{h}^{k}$ (top row), $c_{h}^{k}$ (middle row) and $\phi_{h}^{k}$ (bottom row) displayed at $t_{k}=0.061$ (left) $t_{k}=0.244$ (middle) and $t_{k}=0.427$ (right)
(a) $\delta=0$
(b) $\delta=0.01$
(c) $\delta=0.02$
(d) $\delta=0.05$
(e) $\delta=0.065$
(f) $\delta=0.1$
Hình 2. Comparison of $u_{h}^{k}$ at $t_{k}=0.366$ with $\mu=2$ and different values for the anisotropy strength $\delta$, introduced in (2.15).
(a) isotropic case
(b) $\mu=0$
(c) $\mu=0.5$
(d) $\mu=1.0$
(e) $\mu=1.5$
(f) $\mu=2.0$
Hình 3. Comparison of $u_{h}^{k}$ at $t_{k}=0.122$ with $\delta=0.05$ for different values of $\mu$.
(a) $\beta=0$
(b) $\beta=0.01$
(c) $\beta=0.03$
(d) $\beta=0.1$
Hình 4. Comparison of $u_{h}^{k}$ at $t_{k}=0.305$ with $\delta=0.05$ and $\mu=2$, for solutions to the model with (5.1a) replaced by (5.7) for different values of the noise amplitude $\beta$.
Acknowledgements
AS was partially funded by the European Union (ERC Synergy, NEMESIS, project number 101115663).
Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
OL and VS were partly supported in this work by the Dr Perry James (Jim) Browne Research Centre on Mathematics and its Applications at the University of Sussex, UK.
[…]